Oxycodone, used to relieve moderate to severe pain, is available in 5-mg tablets. A standard dose consists of 5 mg administered at 6-hour intervals. How many tablets will be required for 30 days’ dosage?
Answer #1

ANSWER.

* Calculation of number of tablets required for 30 days.

           Ordered dose = 5 mg at 6 hr intervals or 4 times a day.

                                   = 5 * 4

                                   = 20 mg.

            Available dose = 5 mg/tablet.

    Total number of tablets required per day = 20/5

                                                                       = 4 tablets per day.

Total number of tablets required for 30 days = 4 * 30

                                                                      = 120 tablets.